OpenStudy (anonymous):

1) lim as x --> 0 (x^2 - 2x)/(sin3x)

OpenStudy (anonymous):

I love your Question

OpenStudy (anonymous):

omg are you from tennessee?!

OpenStudy (anonymous):

what do you get to use?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0 } \frac{x^2 -2x}{\sin 3x}\] Apply L'Hospital If it is allowed

OpenStudy (anonymous):

use l'hopital rule ===>

OpenStudy (anonymous):

by which i really mean, can you use l'hopital or do you need something else?

OpenStudy (anonymous):

We're supposed to just use the trig functions

OpenStudy (anonymous):

okay If it is not allowed I still have way to do it

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x^2-2x}{\sin(3x)}=\lim_{x \rightarrow 0}\frac{x(x-2)}{\sin(3x)}=\frac{1}{3} \lim_{x \rightarrow 0}\frac{3x(x-2)}{\sin(3x)}\] \[\frac{1}{3}\lim_{x \rightarrow 0} \frac{3x}{\sin(3x)} \cdot \lim_{x \rightarrow 0}(x-2)\]

myininaya (myininaya):

\[=\frac{1}{3} \cdot 1 \cdot (0-2)\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{x - 2 }{\frac{\sin3x}{x}*\frac{3}{3}}\]

OpenStudy (anonymous):

\[\frac{1}{3}\]

myininaya (myininaya):

\[=\frac{1}{3}(-2)=\frac{-2}{3}\]

OpenStudy (anonymous):

\[\frac{-2}{3}\]Lol typo

OpenStudy (anonymous):

Wow, thank you so much